OpenStudy (anonymous):

If f(x) = 5x - 7, what is f(2)?

OpenStudy (anonymous):

F(2)=5(2)-7=3

OpenStudy (saifoo.khan):

5(2) - 7 10 - 7 3

OpenStudy (anonymous):

f(2) = 5(2) - 7 = 10 - 7 = 3

OpenStudy (anonymous):